The hypotenuse of a right triangle is 34cm. find the length of the two legs, if one leg is 14cm longer than the other.
please answer in complete solution

Answer by nerdybill(7384)

The hypotenuse of a right triangle is 34cm. find the length of the two legs, if one leg is 14cm longer than the other.
.
Let x = short leg
then
x+14 = longer leg
apply Pythagorean theorem, we have:
x^2 + (x+14)^2 = 34^2
x^2 + (x+14)(x+14) = 1156
x^2 + (x^2+28x+196) = 1156
2x^2+28x+196 = 1156
x^2+14x+98 = 578
x^2+14x-480 = 0
(x+16)(x-30) = 0
x = {-16, 30}
we can throw out the negative value (extraneous) leaving:
x = 30 cm (short leg)
.
long leg:
x+14 = 30+14 = 44 cm
.
answer: 30 and 44 cm